A heavy block is suspended from a vertical spring.

The elastic potential energy is stored in the spring is 4 J.

What is the spring constant if the elongation of the spring is 10 cm?


1
Expert's answer
2021-06-11T11:22:19-0400

The elastic potential energy is


"E=\frac 12kx^2,\\\space\\
k=\frac{2E}{x^2}=800\text{ N/m}."